# A Regret Perspective on Online Selective Generation

### v1

@@ G.3 A Proof of Lemma 1
$\textbf{Reg}_{T}$ $=\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\Big{[}a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big{]}-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\left[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\right]$
$\geq\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\Big{[}a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big{]}-\sum_{t=1}^{T}a_{t}(\overline{\tau})-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$ (18)
$=\sum_{t=1}^{T}\left[\mathbbm{E}_{\tau_{t}\sim p_{t}}a_{t}(\tau_{t})-a_{t}(\overline{\tau})\right]+\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$
$\geq-T+\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha),$ (19)
@@
$\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}d_{t}(\tau_{t},\alpha)$ $\leq\frac{T+\textbf{Reg}_{T}}{\lambda}+\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}d_{t}({\tau},\alpha)\leq\frac{T+\textbf{Reg}_{T}}{T^{\nicefrac{{1}}{{4}}}}+\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}d_{t}({\tau},\alpha)$
if $\lambda\geq T^{\nicefrac{{1}}{{4}}}$. Then, by the definition of $d_{t}(\tau_{t},\alpha)$, we have
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)+\alpha\right]$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\nicefrac{{1}}{{4}}}}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha),$
@@
$\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}$ $\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)\right]$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\frac{1}{4}}}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha)-\alpha$
$\leq\frac{1+\textbf{Reg}_{T}/T}{T^{\frac{1}{4}}}$ (20)
@@
Thus, by dividing $1-\textbf{Ineff}_{T}=\frac{1}{T}\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK})$ for both sides of (20), we have
$\textbf{FDR}_{T}=\frac{\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq_{E}\mathbf{y}_{t}\right)}{\sum_{t=1}^{T}\mathbbm{E}_{\tau_{t}\sim p_{t}}\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)}\leq\alpha+\frac{1+\textbf{Reg}_{T}/T}{(1-\textbf{Ineff}_{T})T^{\nicefrac{{1}}{{4}}}}$
@@
G.4 A Proof of Theorem 2
We derive a novel regret bound of our algorithm. The proof technique follows that of the Exp3 regret bound except that we use a novel unbiased estimator for feedback unlocking.
First, we introduce some properties of $\mathcal{{H}}_{t}(\tau_{t})$ in our algorithm. The first trivial property is that $\tau$ is in $\mathcal{{H}}_{t}(\tau)$ itself:
@@ G.4 A Proof of Theorem 2
The second property is that if $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)$, then we have
@@
The third property, similar to the second one, is that if $\tau^{\prime}\in\mathcal{{H}}_{t}^{C}(\tau)$, then we have
$\mathcal{{H}}_{t}^{C}(\tau)=\mathcal{{H}}_{t}(\tau^{\prime}),\text{or equivalently}\mathcal{{H}}_{t}(\tau)=\mathcal{{H}}_{t}^{C}(\tau^{\prime}).$ (23)
@@
The final property is that
@@
Recall that $\ell_{t}(\tau)\in[0,\ell_{\text{max}}]$ and the definition of our loss estimator (8). Then, we observe that this estimator is unbiased as follows:
$\mathbbm{E}_{\tau_{t}\sim p_{t}}\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}({\tau_{t}}))$ $=\sum_{\tau^{\prime}\in\mathcal{{H}}}p_{t}(\tau^{\prime})\left[\frac{\ell_{t}(\tau,\alpha)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))\right]$
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))$
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}({\tau^{\prime}\in\mathcal{{H}}_{t}({\tau})})$ (25)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (26)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (27)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\bar{\tau}\in\mathcal{{H}}_{t}({\tau}))\cdot p_{t}(\bar{\tau})}$ (28)
$=\ell_{t}(\tau,\alpha)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}$
$=\ell_{t}(\tau,\alpha).$ (29)

### v2

@@ H.3 A Proof of Lemma 1
$(1+\lambda)\textbf{Reg}_{T}$ $=\sum_{t=1}^{T}\Big[a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big]-\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}\left[a_{t}(\tau)+\lambda d_{t}(\tau,\alpha)\right]$
$\geq\sum_{t=1}^{T}\Big[a_{t}(\tau_{t})+\lambda d_{t}(\tau_{t},\alpha)\Big]-\sum_{t=1}^{T}a_{t}(\overline{\tau})-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$ (37)
$=\sum_{t=1}^{T}\left[a_{t}(\tau_{t})-a_{t}(\overline{\tau})\right]+\sum_{t=1}^{T}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha)$
$\geq T(\textbf{Ineff}_{T}-1)+\sum_{t=1}^{T}\lambda d_{t}(\tau_{t},\alpha)-\sum_{t=1}^{T}\lambda d_{t}(\overline{\tau},\alpha),$ (38)
@@
$\sum_{t=1}^{T}d_{t}(\tau_{t},\alpha)$ $\leq\frac{T(1-\textbf{Ineff}_{T})+(1+\lambda)\textbf{Reg}_{T}}{\lambda}+\min_{\tau\in\mathcal{{H}}}\sum_{t=1}^{T}d_{t}({\tau},\alpha)$
Then, by the definition of $d_{t}(\tau_{t},\alpha)$, we have
$\frac{1}{T}\sum_{t=1}^{T}$ $\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge e_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)+\alpha\right]$
$\leq\frac{(1-\textbf{Ineff}_{T})+(1+\lambda)\textbf{Reg}_{T}/T}{\lambda}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha),$
@@
$\mathcal{{R}}_{T}^{\textbf{FDR}}$ $=\frac{1}{T}\sum_{t=1}^{T}\left[\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\wedge e_{t}\right)-\alpha\mathbbm{1}\left(\hat{{S}}(\mathbf{x}_{t};\tau_{t})\neq\texttt{IDK}\right)\right]$
$\leq\frac{(1-\textbf{Ineff}_{T})+(1+\lambda)\textbf{Reg}_{T}/T}{\lambda}+\min_{\tau\in\mathcal{{H}}}\frac{1}{T}\sum_{t=1}^{T}d_{t}({\tau},\alpha)-\alpha$
$\leq\frac{(1-\textbf{Ineff}_{T})+{(1+\lambda)\textbf{Reg}_{T}/T}}{\lambda},$ (39)
@@
If we take $\lambda=T^{\nicefrac{{1}}{{4}}}$,
$\mathcal{{R}}_{T}^{\textbf{FDR}}\leq\frac{(1-\textbf{Ineff}_{T})}{T^{\nicefrac{{1}}{{4}}}}+\frac{(1+T^{\nicefrac{{1}}{{4}}})\textbf{Reg}_{T}}{T^{\nicefrac{{5}}{{4}}}},$
@@
Remark.
Finally, we clarify that $\textbf{Ineff}_{T}$ depends on the capability of $G$, the calibration of $f$, and the choice of $\lambda$. In particular, a larger $\lambda$ accelerates convergence but also increases $\textbf{Ineff}_{T}$. Therefore, in practice, choosing a moderately small $\lambda$ can achieve high efficiency while still satisfying the desired FDR under fast convergence.
Nevertheless, if the model or the scoring function $f$ is severely inaccurate, one may heuristically increase $\lambda$ to a large value so that the learner approaches the performance of the best expert that always guarantees the FDR if achievable (Section I for discussion on selection efficiency). In this sense, the exponent $d\in\mathbb{R}^{+}$ in $\lambda=\mathcal{O}(T^{d})$ acts as a hyperparameter of the algorithm.
In our implementation and experiments, we set $\lambda=T^{1/4}$. For empirical trends under different values of $\lambda$, see Figure 26 and 27.
H.4 A Proof of Theorem 1
Here, we derive a novel regret bound of our algorithm, Theorem 4, which is a general case of our main Theorem 1 when the loss range is arbitrarily bounded, i.e., $\in\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$ and the learning rate $\eta$ is non-increasing function in time $t$.
Theorem 4.
Let $\ell_{t}(\cdot)\in[0,\ell_{\text{max}}]$ with the form of (4). For any $T\in\mathbb{N}$ and finite hypotheses $\mathcal{{H}}$, Algorithm 1 provides the following regret bound with probability at least $1-\delta$ if $\eta_{t}=\frac{2\gamma_{t}}{\ell_{\text{max}}}=\sqrt{\frac{\ln|\mathcal{H}|}{\ell_{\text{max}}^{2}T}}$
$\textbf{Reg}_{T}\leq\ell_{\text{max}}\left(4\sqrt{T\ln|\mathcal{H}|}+\left(1+\sqrt{\frac{T}{\ln|\mathcal{{H}}|}}\right)\ln\frac{2}{\delta}\right).$
The proof technique follows that of the Exp3-IX (Neu, 2015) regret bound except that we use a novel loss estimator for feedback unlocking. Note that we abbreviate $\ell_{t}(\tau,\alpha\mid\mathcal{{H}}_{t}(\tau_{t}))$ as $\ell_{t}(\tau\mid\mathcal{{H}}_{t}(\tau_{t}))$ in this proof.
First, we introduce some properties of $\mathcal{{H}}_{t}(\tau_{t})$ in our algorithm.
[…]
(a) Case for $\tau>f_{t}$
[…]
(b) Case for $\tau\leq f_{t}$
Figure 29: Visualization of $\mathcal{{H}}_{t}(\tau)$.
  1. 1.
    (self-inclusion) The first trivial property is that $\tau$ is in $\mathcal{{H}}_{t}(\tau)$ itself:
@@ H.4 A Proof of Theorem 1
  2. 2.
    (partition) The second property is that if $\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)$, then we have
@@
  3. 3.
    (complement) The third property, similar to the second one, is that if $\tau^{\prime}\in\mathcal{{H}}_{t}^{C}(\tau)$, then we have
[…]
@@
  4. 4.
    (swap) The final property is that
@@
With these properties, we can derive the following two useful equalities. Let
$\hat{\ell}_{t}(\tau\mid\mathcal{{H}}(\tau_{t}))=\frac{\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau_{t}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau_{t}})),$
which is our loss estimator without $\gamma_{t}$. Then, the following equality holds:
$\mathbbm{E}_{\tau_{t}\sim p_{t}}\hat{\ell}_{t}(\tau\mid\mathcal{{H}}_{t}({\tau_{t}}))$ $=\sum_{\tau^{\prime}\in\mathcal{{H}}}p_{t}(\tau^{\prime})\left[\frac{\ell_{t}(\tau)}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))\right]$
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}({\tau^{\prime}})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\tau^{\prime}}))$
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}\mathbbm{1}({\tau^{\prime}\in\mathcal{{H}}_{t}({\tau})})$ (44)
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau^{\prime})}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (45)
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\tau\in\mathcal{{H}}_{t}({\bar{\tau}}))\cdot p_{t}(\bar{\tau})}$ (46)
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}\mathbbm{1}(\bar{\tau}\in\mathcal{{H}}_{t}({\tau}))\cdot p_{t}(\bar{\tau})}$ (47)
$=\ell_{t}(\tau)\sum_{\tau^{\prime}\in\mathcal{{H}}_{t}(\tau)}\frac{p_{t}(\tau^{\prime})}{\sum_{\bar{\tau}\in\mathcal{{H}}_{t}(\tau)}p_{t}(\bar{\tau})}$
$=\ell_{t}(\tau),$ (48)